OpenStudy (anonymous):

Part 1: Show the expanded form and simplification of x^6/x^2. Explain in your own words how you can simplify x^6/x^2 without having to write the expanded form. Part 2: Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (nurali):

#1. x^6 / x^2 = (x * x * x * x * x * x) / (x * x) when expanded = (x * x * x * x) after cancelling x terms in the numerator and denominator = x^4 "You can simplify x^6 / x^2 without having to write the expanded form by subtracting the exponent of the numerator by the exponent of like-based factors in the denominator. In this example, the answer is x^(6 - 2) = x^4." #2. (9y^5) / (3y^2) = (3^2 * y^5) / (3^1 * y^2) = 3^(2-1) * y^(5-2) = 3^1 * y^3 = 3y^3